# A Class of De Giorgi Type and Hölder Continuity for Some Problems in Musielak-Orlicz-Sobolev Spaces

### v1

@@ 4.1. Boundedness of Weak Solutions to Problems with Dirichlet Boundary Condition
In this subsection, we focus on proving the boundedness of weak solutions to the Dirichlet problem ($\mathcal{PD}$), as stated in Theorem 1.9. To this end, we need the following lemma concerning the geometric convergence of sequences of numbers will be needed for the De Giorgi iteration arguments below. It can be found in Ho-Sim [44, Lemma 4.3]. The case $\mu_{1}=\mu_{2}$ is contained in Ladyženskaja-Solonnikov-Uralceva [49, Chapter II, Lemma 5.6], see also DiBenedetto [22, Chapter I, Lemma 4.1].
@@
where each $\Omega_{i}:=B_{i}\cap\Omega$ (for $i=1,\ldots,m$) is also a Lipschitz domain (see, for example, Papageorgiou–Winkert [77, Theorem 1.4.86]).
@@
Furthermore, by Propositions 2.11 and 2.16, we have
@@
Let $i\in\{1,\ldots,m\}$. From (4.1), Propositions 2.11, 2.16 and 2.17 for $\Omega=\Omega_{i}$ we have
@@
$\leq C_{2}\left[\|\nabla(u-k_{h+1})^{+}\|_{L^{G(x,t)}(\Omega_{i})}+\|(u-k_{h+1})^{+}\|_{L^{G(x,t)}(\Omega_{i})}\right]^{h^{-}_{1}}.$
Then, again by Propositions 2.11, 2.16 and 2.17 along with (4.1), this leads to
$\int_{\Omega_{i}}H_{1}(x,(u-k_{h+1})^{+}){\rm d}x$ $\leq C_{3}\left(\int_{\Omega_{i}}G(x,|\nabla(u-k_{h+1})^{+}|){\rm d}x+\int_{\Omega_{i}}G(x,(u-k_{h+1})^{+})dx\right)^{\frac{h^{-}}{g^{+}}}$
$\leq C_{3}\left(\int_{A_{k_{h+1}}}G(x,|\nabla u|)dx+\int_{A_{k_{h+1}}}H_{1}(x,(u-k_{h+1})){\rm d}x+|A_{k_{h+1}}|\right)^{\frac{h^{-}_{1}}{g^{+}}}.$
@@
$\int_{A_{k_{h+1}}}G(x,|\nabla u|){\rm d}x\leq C_{6}2^{h\left[\frac{\left(h^{+}\right)^{2}}{g^{-}}+h^{+}\right]}\left(Z_{h-1}^{1+\mu_{1}}+Z_{h-1}^{1+\mu_{2}}\right)\quad\text{for all }h=1,2,\ldots.$
@@
$\int_{\Omega}A(x,u,\nabla u)\cdot\nabla(u-k_{h+1})^{+}dx=\int_{\Omega}B(x,u,\nabla u)(u-k_{h+1})^{+}{\rm d}x,$
@@
$\int_{A_{k_{h+1}}}A(x,u,\nabla u)\cdot\nabla udx=\int_{A_{k_{h+1}}}B(x,u,\nabla u)(u-k_{h+1})dx.$ (4.8)
@@
Since $u\geq u-k_{h+1}>0$, and $u>k_{h+1}\geq 1$ on the set $A_{k_{h+1}}$. Furthermore, the function $t\mapsto H_{1}(x,t)$ is increasing for all $x\in\Omega$. Therefore, by assumption ($\mathcal{A}_{1}$) and Propositions 2.11 and 2.16, we obtain
$\int_{A_{k_{h+1}}}A(x,u,\nabla u)\cdot\nabla u\,\mathrm{d}x$ $\geq a_{4}\int_{A_{k_{h+1}}}G(x,|\nabla u|)\,\mathrm{d}x-a_{5}\int_{A_{k_{h+1}}}H_{1}(x,u)\,\mathrm{d}x-\int_{A_{k_{h+1}}}a_{6}\,\mathrm{d}x$
$\geq a_{4}\int_{A_{k_{h+1}}}G(x,|\nabla u|)\,\mathrm{d}x-\max\{a_{5},a_{6}\}\int_{A_{k_{h+1}}}H_{1}(x,u)\,\mathrm{d}x.$
@@
Then, employing assumption ($\mathcal{B}$), Young’s inequality (2.3), the estimate (2.4), and Propositions 2.11 and 2.16, we find
$\int_{A_{k_{h+1}}}B(x,u,\nabla u)(u-k_{h+1})\,\mathrm{d}x$ $\leq b_{1}\int_{A_{k_{h+1}}}M(x,|\nabla u|)u\,\mathrm{d}x+b_{2}\int_{A_{k_{h+1}}}h_{1}(x,u)u\,\mathrm{d}x+\int_{A_{k_{h+1}}}b_{3}(u-k_{h+1})\,\mathrm{d}x$
$\leq b_{1}\int_{A_{k_{h+1}}}\frac{\varepsilon}{\varepsilon}M(x,|\nabla u|)u\,\mathrm{d}x+\max\{h^{+}_{1}b_{2},b_{3}\}\int_{A_{k_{h+1}}}H_{1}(x,u)\,\mathrm{d}x$
$\leq b_{1}\int_{A_{k_{h+1}}}\widetilde{H_{1}}\left(x,\varepsilon M(x,|\nabla u|)\right)\,\mathrm{d}x+b_{1}\int_{A_{k_{h+1}}}H_{1}\left(x,\frac{1}{\varepsilon}u\right)\,\mathrm{d}x$
@@
$\leq b_{1}\varepsilon^{\frac{h^{+}_{1}}{h^{+}_{1}-1}}\int_{A_{k_{h+1}}}\widetilde{H_{1}}\left(x,M(x,|\nabla u|)\right)\,\mathrm{d}x+b_{1}\varepsilon^{-h^{+}_{1}}\int_{A_{k_{h+1}}}H_{1}(x,u)\,\mathrm{d}x$
@@
$\leq b_{1}\varepsilon^{\frac{h^{+}_{1}}{h^{+}_{1}-1}}\int_{A_{k_{h+1}}}G(x,|\nabla u|)\,\mathrm{d}x+C_{7}\int_{A_{k_{h+1}}}H_{1}(x,u)\,\mathrm{d}x,$
@@
$\int_{A_{k_{h+1}}}B(x,u,\nabla u)(u-k_{h+1})\,\mathrm{d}x\leq\frac{a_{4}}{4}\int_{A_{k_{h+1}}}G(x,|\nabla u|)\,\mathrm{d}x+C_{7}\int_{A_{k_{h+1}}}H_{1}(x,u)\,\mathrm{d}x.$
@@
$\int_{A_{k_{*}}}G(x,|\nabla u|)dx+\int_{A_{k_{*}}}H_{1}(x,u){\rm d}x$ $\leq\min\left\{1,(2\tilde{b}C_{10})^{-\frac{1}{\mu_{1}}}\ \tilde{b}^{-\frac{1}{\mu_{1}^{2}}},\left(2\tilde{b}C_{10}\right)^{-\frac{1}{\mu_{2}}}\ \tilde{b}^{-\frac{1}{\mu_{1}\mu_{2}}-\frac{\mu_{2}-\mu_{1}}{\mu_{2}^{2}}}\right\}.$
@@
$Z_{h}=\int_{A_{k_{h}}}G(x,|\nabla u|)dx+\int_{A_{k_{h}}}H_{1}(x,(u-k_{h}))dx\to 0\quad\text{as }h\to\infty.$
@@ 4.2. Boundedness of Weak Solutions to Problems with Neumann Boundary Condition
$Z_{h}:=\int_{A_{k_{h}}}G(x,|\nabla u|){\rm d}x+\int_{A_{k_{h}}}H_{1}\left(x,(u-k_{h})\right){\rm d}x+\int_{\Gamma_{k_{n}}}H_{2}\left(x,(u-k_{h})\right){\rm d}\sigma,$ (4.17)
@@
Let $i\in\{1,\ldots,m\}$. From Propositions 2.11 and 2.16, inequality (4.16), and Proposition 2.17, for $\Omega=\Omega_{i}$, we have
@@
$\leq C_{3}\left[\|\nabla(u-k_{n+1})^{+}\|_{L^{G(x,t)}(\Omega_{i})}+\|(u-k_{h+1})^{+}\|_{L^{G(x,t)}(\Omega_{i})}\right]^{h_{2}^{-}}.$
Then, by Propositions 2.11, 2.16 and inequality(4.16) it follows
@@
$\leq C_{4}\left(\int_{A_{k_{h+1}}}G(x,|\nabla u|){\rm d}x+\int_{A_{k_{h+1}}}H_{2}(x,(u-k_{h+1})){\rm d}x+|A_{k_{h+1}}|\right)^{\frac{h_{2}^{-}}{g^{+}}}.$
@@
$\int_{A_{k_{h+1}}}G(x,|\nabla u|){\rm d}x\leq C_{7}2^{h\left[\frac{\left(h_{1}^{+}+h_{2}^{+}\right)^{2}}{g^{-}}+g_{*}^{+}\right]}\left(Z_{h-1}^{1+\mu_{1}}+Z_{h-1}^{1+\mu_{2}}\right)\quad\text{for all }h=1,2,3,\ldots,$
@@
$\int_{\Omega}A(x,u,\nabla u)\cdot\nabla\varphi{\rm d}x=\int_{\Omega}B(x,u,\nabla u)\varphi{\rm d}x+\int_{\partial\Omega}C(x,u)\varphi{\rm d}\sigma,$
@@
$\begin{split}&\int_{A_{k_{h+1}}}A(x,u,\nabla u)\cdot\nabla u{\rm d}x\\ &=\int_{A_{k_{h+1}}}B(x,u,\nabla u)(u-k_{h+1}){\rm d}x+\int_{\Gamma_{k_{h+1}}}C(x,u)(u-k_{h+1}){\rm d}\sigma.\end{split}$ (4.24)
As done in the proof of Theorem 1.9, we apply assumptions ($\mathcal{A}_{1}$), ($\mathcal{B}$) and ($\mathcal{C}$), as well as Young’s inequality (2.3), (2.4) and Propositions 2.11 and 2.16, and we take $\varepsilon\in(0,1]$ such that
@@
$\begin{split}&\int_{A_{k_{h+1}}}A(x,u,\nabla u)\cdot\nabla u{\rm d}x\\ &\geq a_{4}\int_{A_{k_{h+1}}}G(x,|\nabla u|){\rm d}x-\max\{a_{5},a_{6}\}\int_{A_{k_{h+1}}}H_{1}(x,u){\rm d}x,\end{split}$ (4.25)
@@
$\begin{split}&\int_{A_{k_{h+1}}}B(x,u,\nabla u)(u-k_{h+1}){\rm d}x\\ &\leq b_{1}\int_{A_{k_{h+1}}}M(x,|\nabla u|)(u-k_{h+1}){\rm d}x+b_{2}\int_{A_{k_{h+1}}}h_{1}(x,u)(u-k_{h+1}){\rm d}x+b_{3}\int_{A_{k_{h+1}}}(u-k_{h+1})dx\\ &\leq b_{1}\int_{A_{k_{h+1}}}M(x,|\nabla u|)u{\rm d}x+b_{2}\int_{A_{k_{h+1}}}h_{1}(x,u)u{\rm d}x+b_{3}\int_{A_{k_{h+1}}}u(x)dx\\ &\leq b_{1}\int_{A_{k_{h+1}}}M(x,|\nabla u|)u{\rm d}x+b_{2}h_{1}^{+}\int_{A_{k_{h+1}}}H_{1}(x,u){\rm d}x+b_{3}\int_{A_{k_{h+1}}}H_{1}(x,u)dx\\ &\leq b_{1}\int_{A_{k_{h+1}}}\frac{\varepsilon}{\varepsilon}M(x,|\nabla u|)u{\rm d}x+\max\{h_{1}^{+}b_{2},b_{3}\}\int_{A_{k_{h+1}}}H_{1}(x,u){\rm d}x\\ &\leq b_{1}\int_{A_{k_{h+1}}}\widetilde{H_{1}}\left(x,\varepsilon M(x,|\nabla u|)\right){\rm d}x+b_{1}\int_{A_{k_{h+1}}}H_{1}\left(x,\frac{1}{\varepsilon}u\right){\rm d}x\\ &+\max\{h_{1}^{+}b_{2},b_{3}\}\int_{A_{k_{h+1}}}H_{1}(x,u){\rm d}x\\ &\leq b_{1}\varepsilon^{\frac{h_{1}^{+}}{h_{1}^{+}-1}}\int_{A_{k_{h+1}}}\widetilde{H_{1}}\left(x,M(x,|\nabla u|)\right){\rm d}x+b_{1}\varepsilon^{-h_{1}^{+}}\int_{A_{k_{h+1}}}H_{1}\left(x,u\right){\rm d}x\\ &+\max\{h_{1}^{+}b_{2},b_{3}\}\int_{A_{k_{h+1}}}H_{1}(x,u){\rm d}x\\ &\leq b_{1}\varepsilon^{\frac{h_{1}^{+}}{h_{1}^{+}-1}}\int_{A_{k_{h+1}}}G(x,|\nabla u|){\rm d}x+\max\{h_{1}^{+}b_{2},b_{3},b_{1}\varepsilon^{-h_{1}^{+}}\}\int_{A_{k_{h+1}}}H_{1}(x,u){\rm d}x\\ &=\frac{a_{4}}{4}\int_{A_{k_{h+1}}}G(x,|\nabla u|){\rm d}x+C_{7}\int_{A_{k_{h+1}}}H_{1}(x,u){\rm d}x.\end{split}$ (4.26)
@@
$\int_{A_{k_{h+1}}}G(x,|\nabla u|){\rm d}x\leq C_{9}2^{h(h_{1}^{+}+h_{2}^{+})}\left[\int_{A_{k_{h}}}H_{1}(x,(u-k_{h})){\rm d}x+\int_{\Gamma_{k_{h}}}H_{2}(x,(u-k_{h})){\rm d}\sigma\right].$

### v2

@@ 4.1. Boundedness of Weak Solutions to Problems with Dirichlet Boundary Condition
In this subsection, we focus on proving the boundedness of weak solutions to the Dirichlet problem ($\mathcal{PD}$), as stated in Theorem 1.9. To this end, we need the following lemma concerning the geometric convergence of sequences of numbers will be needed for the De Giorgi iteration arguments below. It can be found in Ho-Sim [42, Lemma 4.3]. The case $\mu_{1}=\mu_{2}$ is contained in Ladyženskaja-Solonnikov-Uralceva [46, Chapter II, Lemma 5.6], see also DiBenedetto [19, Chapter I, Lemma 4.1].
@@
where each $\Omega_{i}:=B_{i}\cap\Omega$ (for $i=1,\ldots,m$) is also a Lipschitz domain (see, for example, Papageorgiou–Winkert [74, Theorem 1.4.86]).
@@
Furthermore, by Propositions 2.12 and 2.17, we have
@@
Let $i\in\{1,\ldots,m\}$. From (4.1), Propositions 2.12, 2.17 and 2.18 for $\Omega=\Omega_{i}$ we have
@@
$\leq C_{2}\left[\|D(u-k_{h+1})^{+}\|_{L^{G(x,t)}(\Omega_{i})}+\|(u-k_{h+1})^{+}\|_{L^{G(x,t)}(\Omega_{i})}\right]^{h^{-}_{1}}.$
Then, again by Propositions 2.12, 2.17 and 2.18 along with (4.1), this leads to
$\int_{\Omega_{i}}H_{1}(x,(u-k_{h+1})^{+}){\rm d}x$ $\leq C_{3}\left(\int_{\Omega_{i}}G(x,|D(u-k_{h+1})^{+}|){\rm d}x+\int_{\Omega_{i}}G(x,(u-k_{h+1})^{+})dx\right)^{\frac{h^{-}}{g^{+}}}$
$\leq C_{3}\left(\int_{A_{k_{h+1}}}G(x,|Du|)dx+\int_{A_{k_{h+1}}}H_{1}(x,(u-k_{h+1})){\rm d}x+|A_{k_{h+1}}|\right)^{\frac{h^{-}_{1}}{g^{+}}}.$
@@
$\int_{A_{k_{h+1}}}G(x,|Du|){\rm d}x\leq C_{6}2^{h\left[\frac{\left(h^{+}\right)^{2}}{g^{-}}+h^{+}\right]}\left(Z_{h-1}^{1+\mu_{1}}+Z_{h-1}^{1+\mu_{2}}\right)\quad\text{for all }h=1,2,\ldots.$
@@
$\int_{\Omega}A(x,u,Du)\cdot D(u-k_{h+1})^{+}dx=\int_{\Omega}B(x,u,Du)(u-k_{h+1})^{+}{\rm d}x,$
@@
$\int_{A_{k_{h+1}}}A(x,u,Du)\cdot Dudx=\int_{A_{k_{h+1}}}B(x,u,Du)(u-k_{h+1})dx.$ (4.8)
@@
Since $u\geq u-k_{h+1}>0$, and $u>k_{h+1}\geq 1$ on the set $A_{k_{h+1}}$. Furthermore, the function $t\mapsto H_{1}(x,t)$ is increasing for all $x\in\Omega$. Therefore, by assumption ($\mathcal{A}_{1}$) and Propositions 2.12 and 2.17, we obtain
$\int_{A_{k_{h+1}}}A(x,u,Du)\cdot Du\,\mathrm{d}x$ $\geq a_{4}\int_{A_{k_{h+1}}}G(x,|Du|)\,\mathrm{d}x-a_{5}\int_{A_{k_{h+1}}}H_{1}(x,u)\,\mathrm{d}x-\int_{A_{k_{h+1}}}a_{6}\,\mathrm{d}x$
$\geq a_{4}\int_{A_{k_{h+1}}}G(x,|Du|)\,\mathrm{d}x-\max\{a_{5},a_{6}\}\int_{A_{k_{h+1}}}H_{1}(x,u)\,\mathrm{d}x.$
@@
Then, employing assumption ($\mathcal{B}$), Young’s inequality (2.3), the estimate (2.4), and Propositions 2.12 and 2.17, we find
$\int_{A_{k_{h+1}}}B(x,u,Du)(u-k_{h+1})\,\mathrm{d}x$ $\leq b_{1}\int_{A_{k_{h+1}}}M(x,|Du|)u\,\mathrm{d}x+b_{2}\int_{A_{k_{h+1}}}h_{1}(x,u)u\,\mathrm{d}x+\int_{A_{k_{h+1}}}b_{3}(u-k_{h+1})\,\mathrm{d}x$
$\leq b_{1}\int_{A_{k_{h+1}}}\frac{\varepsilon}{\varepsilon}M(x,|Du|)u\,\mathrm{d}x+\max\{h^{+}_{1}b_{2},b_{3}\}\int_{A_{k_{h+1}}}H_{1}(x,u)\,\mathrm{d}x$
$\leq b_{1}\int_{A_{k_{h+1}}}\widetilde{H_{1}}\left(x,\varepsilon M(x,|Du|)\right)\,\mathrm{d}x+b_{1}\int_{A_{k_{h+1}}}H_{1}\left(x,\frac{1}{\varepsilon}u\right)\,\mathrm{d}x$
@@
$\leq b_{1}\varepsilon^{\frac{h^{+}_{1}}{h^{+}_{1}-1}}\int_{A_{k_{h+1}}}\widetilde{H_{1}}\left(x,M(x,|Du|)\right)\,\mathrm{d}x+b_{1}\varepsilon^{-h^{+}_{1}}\int_{A_{k_{h+1}}}H_{1}(x,u)\,\mathrm{d}x$
@@
$\leq b_{1}\varepsilon^{\frac{h^{+}_{1}}{h^{+}_{1}-1}}\int_{A_{k_{h+1}}}G(x,|Du|)\,\mathrm{d}x+C_{7}\int_{A_{k_{h+1}}}H_{1}(x,u)\,\mathrm{d}x,$
@@
$\int_{A_{k_{h+1}}}B(x,u,Du)(u-k_{h+1})\,\mathrm{d}x\leq\frac{a_{4}}{4}\int_{A_{k_{h+1}}}G(x,|Du|)\,\mathrm{d}x+C_{7}\int_{A_{k_{h+1}}}H_{1}(x,u)\,\mathrm{d}x.$
@@
$\int_{A_{k_{*}}}G(x,|Du|)dx+\int_{A_{k_{*}}}H_{1}(x,u){\rm d}x$ $\leq\min\left\{1,(2\tilde{b}C_{10})^{-\frac{1}{\mu_{1}}}\ \tilde{b}^{-\frac{1}{\mu_{1}^{2}}},\left(2\tilde{b}C_{10}\right)^{-\frac{1}{\mu_{2}}}\ \tilde{b}^{-\frac{1}{\mu_{1}\mu_{2}}-\frac{\mu_{2}-\mu_{1}}{\mu_{2}^{2}}}\right\}.$
@@
$Z_{h}=\int_{A_{k_{h}}}G(x,|Du|)dx+\int_{A_{k_{h}}}H_{1}(x,(u-k_{h}))dx\to 0\quad\text{as }h\to\infty.$
@@ 4.2. Boundedness of Weak Solutions to Problems with Neumann Boundary Condition
$Z_{h}:=\int_{A_{k_{h}}}G(x,|Du|){\rm d}x+\int_{A_{k_{h}}}H_{1}\left(x,(u-k_{h})\right){\rm d}x+\int_{\Gamma_{k_{n}}}H_{2}\left(x,(u-k_{h})\right){\rm d}\sigma,$ (4.17)
@@
Let $i\in\{1,\ldots,m\}$. From Propositions 2.12 and 2.17, inequality (4.16), and Proposition 2.18, for $\Omega=\Omega_{i}$, we have
@@
$\leq C_{3}\left[\|D(u-k_{n+1})^{+}\|_{L^{G(x,t)}(\Omega_{i})}+\|(u-k_{h+1})^{+}\|_{L^{G(x,t)}(\Omega_{i})}\right]^{h_{2}^{-}}.$
Then, by Propositions 2.12, 2.17 and inequality(4.16) it follows
@@
$\leq C_{4}\left(\int_{A_{k_{h+1}}}G(x,|Du|){\rm d}x+\int_{A_{k_{h+1}}}H_{2}(x,(u-k_{h+1})){\rm d}x+|A_{k_{h+1}}|\right)^{\frac{h_{2}^{-}}{g^{+}}}.$
@@
$\int_{A_{k_{h+1}}}G(x,|Du|){\rm d}x\leq C_{7}2^{h\left[\frac{\left(h_{1}^{+}+h_{2}^{+}\right)^{2}}{g^{-}}+g_{*}^{+}\right]}\left(Z_{h-1}^{1+\mu_{1}}+Z_{h-1}^{1+\mu_{2}}\right)\quad\text{for all }h=1,2,3,\ldots,$
@@
$\int_{\Omega}A(x,u,Du)\cdot\nabla\varphi{\rm d}x=\int_{\Omega}B(x,u,Du)\varphi{\rm d}x+\int_{\partial\Omega}C(x,u)\varphi{\rm d}\sigma,$
@@
$\begin{split}&\int_{A_{k_{h+1}}}A(x,u,Du)\cdot Du{\rm d}x\\ &=\int_{A_{k_{h+1}}}B(x,u,Du)(u-k_{h+1}){\rm d}x+\int_{\Gamma_{k_{h+1}}}C(x,u)(u-k_{h+1}){\rm d}\sigma.\end{split}$ (4.24)
As done in the proof of Theorem 1.9, we apply assumptions ($\mathcal{A}_{1}$), ($\mathcal{B}$) and ($\mathcal{C}$), as well as Young’s inequality (2.3), (2.4) and Propositions 2.12 and 2.17, and we take $\varepsilon\in(0,1]$ such that
@@
$\begin{split}&\int_{A_{k_{h+1}}}A(x,u,Du)\cdot Du{\rm d}x\\ &\geq a_{4}\int_{A_{k_{h+1}}}G(x,|Du|){\rm d}x-\max\{a_{5},a_{6}\}\int_{A_{k_{h+1}}}H_{1}(x,u){\rm d}x,\end{split}$ (4.25)
@@
$\begin{split}&\int_{A_{k_{h+1}}}B(x,u,Du)(u-k_{h+1}){\rm d}x\\ &\leq b_{1}\int_{A_{k_{h+1}}}M(x,|Du|)(u-k_{h+1}){\rm d}x+b_{2}\int_{A_{k_{h+1}}}h_{1}(x,u)(u-k_{h+1}){\rm d}x+b_{3}\int_{A_{k_{h+1}}}(u-k_{h+1})dx\\ &\leq b_{1}\int_{A_{k_{h+1}}}M(x,|Du|)u{\rm d}x+b_{2}\int_{A_{k_{h+1}}}h_{1}(x,u)u{\rm d}x+b_{3}\int_{A_{k_{h+1}}}u(x)dx\\ &\leq b_{1}\int_{A_{k_{h+1}}}M(x,|Du|)u{\rm d}x+b_{2}h_{1}^{+}\int_{A_{k_{h+1}}}H_{1}(x,u){\rm d}x+b_{3}\int_{A_{k_{h+1}}}H_{1}(x,u)dx\\ &\leq b_{1}\int_{A_{k_{h+1}}}\frac{\varepsilon}{\varepsilon}M(x,|Du|)u{\rm d}x+\max\{h_{1}^{+}b_{2},b_{3}\}\int_{A_{k_{h+1}}}H_{1}(x,u){\rm d}x\\ &\leq b_{1}\int_{A_{k_{h+1}}}\widetilde{H_{1}}\left(x,\varepsilon M(x,|Du|)\right){\rm d}x+b_{1}\int_{A_{k_{h+1}}}H_{1}\left(x,\frac{1}{\varepsilon}u\right){\rm d}x\\ &+\max\{h_{1}^{+}b_{2},b_{3}\}\int_{A_{k_{h+1}}}H_{1}(x,u){\rm d}x\\ &\leq b_{1}\varepsilon^{\frac{h_{1}^{+}}{h_{1}^{+}-1}}\int_{A_{k_{h+1}}}\widetilde{H_{1}}\left(x,M(x,|Du|)\right){\rm d}x+b_{1}\varepsilon^{-h_{1}^{+}}\int_{A_{k_{h+1}}}H_{1}\left(x,u\right){\rm d}x\\ &+\max\{h_{1}^{+}b_{2},b_{3}\}\int_{A_{k_{h+1}}}H_{1}(x,u){\rm d}x\\ &\leq b_{1}\varepsilon^{\frac{h_{1}^{+}}{h_{1}^{+}-1}}\int_{A_{k_{h+1}}}G(x,|Du|){\rm d}x+\max\{h_{1}^{+}b_{2},b_{3},b_{1}\varepsilon^{-h_{1}^{+}}\}\int_{A_{k_{h+1}}}H_{1}(x,u){\rm d}x\\ &=\frac{a_{4}}{4}\int_{A_{k_{h+1}}}G(x,|Du|){\rm d}x+C_{7}\int_{A_{k_{h+1}}}H_{1}(x,u){\rm d}x.\end{split}$ (4.26)
@@
$\int_{A_{k_{h+1}}}G(x,|Du|){\rm d}x\leq C_{9}2^{h(h_{1}^{+}+h_{2}^{+})}\left[\int_{A_{k_{h}}}H_{1}(x,(u-k_{h})){\rm d}x+\int_{\Gamma_{k_{h}}}H_{2}(x,(u-k_{h})){\rm d}\sigma\right].$